# Im2SurfTex: Surface Texture Generation via Neural Backprojection of Multi-View Images

### v1

@@ Abstract
We present Im2SurfTex, a method that generates textures for input 3D shapes by learning to aggregate multi-view image outputs produced by 2D image diffusion models onto the shapes’ texture space. Unlike existing texture generation techniques that use ad hoc backprojection and averaging schemes to blend multiview images into textures, often resulting in texture seams and artifacts, our approach employs a trained, feedforward neural module to boost texture coherency. The key ingredient of our module is to leverage neural attention and appropriate positional encodings of image pixels based on their corresponding 3D point positions, normals, and surface-aware coordinates as encoded in geodesic distances within surface patches. These encodings capture texture correlations between neighboring surface points, ensuring better texture continuity. Experimental results show that our module improves texture quality, achieving superior performance in high-resolution texture generation.
@@ 1 Introduction
Figure 1: A gallery of 3D shapes from various categories, textured by  Im2SurfTex. Our method is able to produce high-resolution, seamless, and visually coherent textures.
@@
Our approach, named  Im2SurfTex, addresses the above limitations with the introduction of a novel feedforward module that can be easily integrated to existing texture generation approaches. The module is trained to combine color information from multiple viewpoints to textures through a cross-attention mechanism, where for each surface point, several candidate image neighborhoods across different views are examined and combined back in texture space depending on local surface geometry, as encoded in 3D positions, normals and geodesic distances within these patches. In this manner, the attention mechanism captures local context based on surface (geodesic) proximity rather than relying solely on 3D Euclidean proximity that might correlate surface region textures far from each other in geodesic sense. Our module operates in a feed-forward manner, yielding coherent textures efficiently, within a few seconds, without requiring any slow optimization procedures. Our experiments indicate significant improvements in generated texture quality, measured by different scores, including FID [HRU17], KID [BSAG18], CLIP [SBV22] metrics, when compared to alternatives.
@@
  • a cross-attention mechanism that learns how to wrap generated images from different views onto a single, coherent surface texture map in a feed-forward manner.
@@ 2 Related Work
To address these issues, other methods such as TexFusion [CKF23] leverage latent diffusion to interlace diffusion and back-projection steps, producing 3D-aware latent images that are subsequently decoded and merged into a texture map. Similarly, SyncMVD [LXLW24] employs a latent texture map where all views are encoded at each denoising step, further enhancing consistency in geometry and appearance. Still, both methods rely on ad hoc heuristics for aggregating texture information from different views, such as mere averaging or using the most front-facing view information for each texel. In contrast, our approach learns to aggregate color information from multiple views based on both geometry and texture information, enabling the generation of a coherent and seamless surface texture map.
A more recent approach, Paint3D [ZCQ24], achieves impressive texture generation results, by adopting a two-stage texture generation strategy. In the first stage, a coarse texture is created by backprojecting views to texture space via the heuristic of using the most front-facing view information for each texel, as in previous methods. The second stage involves a refinement and inpainting process that utilizes a diffusion model in UV texture space, conditioned on a UV position map encoding the 3D adjacency information of texels. While this method directly encodes 3D geometric information into the texture map, it can still result in misaligned textures due to its employed heuristic during its coarse stage. The subsequent texture refinement steps often fail to fix the artifacts of the coarse stage, as demonstrated in our experiments. Another related method, MatAtlas [CDG24], incorporates a three-step denoising process with sequential operations and line conditions to preserve geometry and style consistency. However, MatAtlas employs an averaging heuristic for blending the final texture from the generated views, leading to inconsistencies or overly smooth surfaces, as also demonstrated in our experiments.
@@ 3.5 Training
3.6 Implementation details
During inference, our approach follows either backbone described in Section 3.1, yet incorporating the learned backprojection module instead of their heuristic backprojection. Since some texels remain unfilled after backprojection, they are subsequently inpainted and refined using pretrained Stable Diffusion 1.5 and the Paint3D’s 3D-aware ControlNet module. The final output textures have a resolution of $1024\times 1024$. The entire texturing process takes about a minute on a single NVIDIA A6000 GPU to texture an input shape. We also refer readers to our project page with source code for more details. [^1] Project page (with code): ygeorg01.github.io/Im2SurfTex
@@ 4 Evaluation
We evaluate  Im2SurfTex on text-to-texture generation both quantitatively and qualitatively. In the following sections, we explain the experimental setup for evaluating our approach, including the evaluation dataset and metrics (Section 4.1). We then compare  Im2SurfTex  against competing text-to-texture generation methods (Section 4.2). We also analyze the impact of pixel neighborhood sizes and geodesic distances on performance in an ablation study (Section 4.3).
@@ 4.1 Experimental setup
We evaluate  Im2SurfTex on the test split provided by Text2Tex [CSL23]. The split includes $410$ textured meshes from Objaverse [DSS23] across $225$ categories. All competing methods are trained on the same training split, as discussed in Section 3.5, and evaluated on the same above test Objaverse split.
@@ 4.2 Comparisons
Our main finding – replacing traditional backprojection with our neural module – is numerically examined in Table 1. Our neural module improves both the original Paint3D backbone as well as our implemented MatAtlas backbone. The improvements are consistent across all three evaluation metrics (FID, KID, and CLIP score). Our neural backprojection improves the FID distance by $6.1\%$ for the Paint3D backbone, and $6.9\%$ for the MatAtlas backbone. The improvements are more prominent in terms of the KID score ($19.1\%$ relative reduction for the Paint3D backbone, and $29.2\%$ relative reduction for the MatAtlas backbone). In terms of CLIP score, all methods seem to generate images that are similarly aligned with the text prompt, yet our module still maintains a small edge over traditional backprojection.
@@
In Table 2, we include quantitative comparisons of the best variant of our method (based on the MatAtlas backbone) with other state-of-the-art models for text-to-texture generation. According to all the evaluation metrics, our method provides the best performance in terms of FID & KID distances as well as CLIP score.
@@
Table 2: Comparisons with other text-to-texture methods
@@ 4.3 Ablation
Table 4 provides another ablation where we examine whether using the geodesic distance as additional feature in the positional encodings of the pixels helps. We observe rather minor improvements in terms of the numerical scores. We suspect that the small differences are due to the fact that the improvements happen only in small image regions for the shapes of our dataset, where the surface changes rapidly (e.g., folds, handles, high curvature regions), as shown in Figure 5. These small regions seem to have a relatively small effect on the established image quality metrics. Figure 5 demonstrates that adding geodesic distances as features in our module leads to fewer texture artifacts and diminished color bleeding in these regions e.g., see the color bleeding between the bed mattress and wooden frame, or the green leaf and the apple when geodedic distances are not used.
@@
Geodesic Distances FID $\downarrow$ KID $\downarrow$ CLIPscore $\uparrow$
- 27.35 2.15 $\pm$ 0.2 29.61
27.34 2.12 $\pm$ 0.3 29.63

### v2

@@ Abstract
We present Im2SurfTex, a method that generates textures for input 3D shapes by learning to aggregate multi-view image outputs produced by 2D image diffusion models onto the shapes’ texture space. Unlike existing texture generation techniques that use ad hoc backprojection and averaging schemes to blend multiview images into textures, often resulting in texture seams and artifacts, our approach employs a trained neural module to boost texture coherency. The key ingredient of our module is to leverage neural attention and appropriate positional encodings of image pixels based on their corresponding 3D point positions, normals, and surface-aware coordinates as encoded in geodesic distances within surface patches. These encodings capture texture correlations between neighboring surface points, ensuring better texture continuity. Experimental results show that our module improves texture quality, achieving superior performance in high-resolution texture generation.
@@ 1 Introduction
Figure 1: A gallery of 3D shapes from various categories, textured by  Im2SurfTex.
@@
Our approach, named  Im2SurfTex, addresses the above limitations with the introduction of a novel, optimization-free module that can be easily integrated to existing texture generation approaches. The module is trained to combine color information from multiple viewpoints to textures through a cross-attention mechanism, where for each surface point, several candidate image neighborhoods across different views are examined and combined back in texture space depending on local surface geometry, as encoded in 3D positions, normals and geodesic distances within these patches. In this manner, the attention mechanism captures local context based on surface (geodesic) proximity rather than relying solely on 3D Euclidean proximity that might correlate surface region textures far from each other in geodesic sense. Our module yields coherent textures efficiently, without requiring any slow optimization procedures. Our experiments indicate significant improvements in generated texture quality, measured by different scores, including FID [HRU17], KID [BSAG18], CLIP [SBV22] metrics, when compared to alternatives.
@@
  • a cross-attention mechanism that learns how to wrap generated images from different views onto a single, coherent surface texture map.
@@ 2 Related Work
To address these issues, other methods such as TexFusion [CKF23] leverage latent diffusion to interlace diffusion and back-projection steps, producing 3D-aware latent images that are subsequently decoded and merged into a texture map. Similarly, SyncMVD [LXLW24] employs a latent texture map where all views are encoded at each denoising step, further enhancing consistency in geometry and appearance. TexGen [HGZ24] introduced a multi-view sampling and resampling framework that updates a UV texture map iteratively during denoising, aiming to reduce view discrepancies. Still, these methods rely on ad hoc blending masks or heuristics for aggregating texture information from different views, such as mere averaging or using the most front-facing view information for each texel. In contrast, our approach learns to aggregate color information from multiple views based on both geometry and texture information, promoting the generation of more coherent and seamless surface texture maps.
A more recent approach, Paint3D [ZCQ24], achieves impressive texture generation results, by adopting a two-stage texture generation strategy. In the first stage, a coarse texture is created by backprojecting views to texture space via the heuristic of using the most front-facing view information for each texel, as in previous methods. The second stage involves a refinement and inpainting process that utilizes a diffusion model in UV texture space, conditioned on a UV position map encoding the 3D adjacency information of texels. While this method directly encodes 3D geometric information into the texture map, it can still result in misaligned textures due to its employed heuristic during its coarse stage. The subsequent texture refinement steps often fail to fix the artifacts of the coarse stage, as demonstrated in our experiments. Another related method, MatAtlas [CDG24], incorporates a three-step denoising process with sequential operations and line conditions to preserve geometry and style consistency. However, MatAtlas employs an averaging heuristic for blending the final texture from the generated views, leading to inconsistencies or overly smooth surfaces, as also demonstrated in our experiments. TEXGen [YYG24] takes a different approach by directly training a large-scale diffusion model in the UV texture space, and integrating convolution operations in UV space with 3D-aware attention layers in their denoising network to achieve high-resolution texture synthesis. However, their method still faces challenges in maintaining cross-view consistency since the generated textures are conditioned on single-view images which are merely backprojected to the UV space to derive the initial partial texture maps used in their diffusion.
@@
Expanding beyond single object texture generation, InstanceTex [YGC24] focuses on texture generation for 3D scenes, employing a local synchronized multi-view diffusion strategy to improve local texture consistency across multiple objects. 3D Paintbrush [DLAH24] specializes in localized stylization of single objects, using cascaded score distillation to refine textures within specific object regions. These approaches differ from our method in scope: InstanceTex is tailored for stylistic consistency in large environments, while 3D Paintbrush targets localized edits.
@@ 3.5 Training
3.6 Implementation details
During inference, our approach follows either backbone described in Section 3.1, yet incorporating the learned backprojection module instead of their heuristic backprojection. Since some texels remain unfilled after backprojection, they are subsequently inpainted and refined using pretrained Stable Diffusion 1.5 and the Paint3D’s 3D-aware ControlNet module. The final output textures have a resolution of $1024\times 1024$. The entire texturing process takes one to two minutes on a single NVIDIA A6000 GPU to texture an input shape. Each iteration of view generation and neural backprojection takes approximately $35$ seconds, while the inpainting and high-definition (HD) modules require around $20$ seconds each. For preprocessing, our approach employs a one-time procedure to compute geodesic information metadata, which takes approximately $30$ minutes when processing a new object for the first time. This part can be significantly accelerated with more efficient techniques for computation of geodesics [CWW13, ZHA23]. We also refer readers to our project page with source code for more details.  [^1] Project page (with code): ygeorg01.github.io/Im2SurfTex
@@ 4 Evaluation
We evaluate  Im2SurfTex on text-to-texture generation both quantitatively and qualitatively. In the following sections, we explain the experimental setup for evaluating our approach, including the evaluation dataset and metrics (Section 4.1). We then compare  Im2SurfTex  against competing text-to-texture generation methods (Section 4.2). We also analyze the impact of pixel neighborhood sizes, geodesic distances and number of input views on performance in an ablation study (Section 4.3).
@@ 4.1 Experimental setup
We evaluate  Im2SurfTex on the test split provided by Text2Tex [CSL23]. The split includes $410$ textured meshes from Objaverse [DSS23] across $225$ categories. All competing methods are trained on the same training split, as discussed in Section 3.5, and evaluated on the same above test Objaverse split. We also note that all competing methods use the same UV maps and surface parametrization.
@@ 4.2 Comparisons
Our main finding – replacing traditional backprojection with our neural module – is numerically examined in Table 1. Our neural module improves both the original Paint3D backbone as well as our implemented MatAtlas backbone. The improvements are consistent across all three evaluation metrics (FID, KID, and CLIP score). Our neural backprojection improves the FID distance by $6.1\%$ for the Paint3D backbone, and $6.9\%$ for the MatAtlas backbone. The improvements are more prominent in terms of the KID score ($19.1\%$ relative reduction for the Paint3D backbone, and $29.2\%$ relative reduction for the MatAtlas backbone). The KID score is more sensitive to fine-grained texture variations due to the use of Maximum Mean Discrepancy (MMD) with a polynomial kernel when comparing distributions. As a result, when a model improvement primarily reduces local inconsistencies – such as texture artifacts and fine details, KID tends to exhibit a more substantial improvement than FID. In terms of CLIP score, all methods seem to generate images that are similarly aligned with the text prompt, yet our module still maintains a small edge over traditional backprojection.
@@
In Table 2, we include quantitative comparisons of the best variant of our method (based on the MatAtlas backbone) with other state-of-the-art models for text-to-texture generation. Here we also include a comparison with the recent method of TEXGen[YYG24]. According to all the evaluation metrics, our method provides the best performance in terms of FID & KID distances as well as CLIP score.
In Figure 5, we show qualiitative comparisons with TEXGen’s released implementation [YYG24]. We observe that TEXGen often leads to global texture inconsistencies on the output shapes, while our method is more view-consistent.
@@
Table 2: Comparisons with other text-to-texture methods.
Figure 5: Comparison between  Im2SurfTex and TEXGen [YYG24]. We show two different views of the textured objects. Our method produces more coherent and view-consistent textures.
@@ 4.3 Ablation
Table 4 provides another ablation where we compare using absolute versus relative coordinates in the positional encodings of the Eq. 4, and also examine whether using geodesic distance as additional feature in the positional encodings helps. Relative coordinates enhance performance compared to absolute coordinates, as they provide a more effective encoding for processing the local interactions between neighboring points, regardless of their actual 3D locations. With repect to the use of geodesic distances, we observe rather minor improvements in terms of the numerical scores. We suspect that the small differences are due to the fact that the improvements happen only in small image regions for the shapes of our dataset, where the surface changes rapidly (e.g., folds, handles, high curvature regions), as shown in Figure 7. These small regions seem to have a relatively small effect on the established image quality metrics. Figure 7 demonstrates that adding geodesic distances as features in our module leads to fewer texture artifacts and diminished color bleeding in these regions e.g., see the color bleeding between the bed mattress and wooden frame, or the green leaf and the apple when geodedic distances are not used.
Figure 6 demonstrates a visual comparison between reference textured meshes from our dataset, and reconstructed textures by our method, when we pass as input the rendered images from the original textures. This comparison aims to show whether our method causes any significant color shifting or bleeding while aggregating information from different views. We see that demonstrating our method does not introduce any such discrepancies during neural backprojection.
Figure 6: Our neural backprojection can closely reconstruct challenging textures of target objects in our dataset without causing noticeable color shifting or discrepancies between target and decoded textures.
@@
Rel Coords. Geod. Distances FID $\downarrow$ KID $\downarrow$ CLIPscore $\uparrow$ - - 27.86 2.32 $\pm$ 0.2 29.62 - 27.35 2.15 $\pm$ 0.2 29.61 27.34 2.12 $\pm$ 0.3 29.63
@@
# of Views Method FID $\downarrow$ KID $\downarrow$ CLIPscore $\uparrow$ 4 Paint3D 29.41 2.71 $\pm$ 0.3 29.40 Im2SurfTex 28.51 2.42 $\pm$ 0.3 29.62 6 Paint3D 29.13 2.62 $\pm$ 0.2 29.45 Im2SurfTex 27.34 2.12 $\pm$ 0.2 29.63 8 Paint3D 29.20 2.75 $\pm$ 0.3 29.48 Im2SurfTex 27.86 2.32 $\pm$ 0.2 29.62
Table 5: Ablation study results wrt using different number of views.
Table 5 presents the impact of using different number of views on our evaluation metrics for both Paint3D and our method (using the Paint3D backbone). Increasing the number of views from $4$ to $6$ views results in improvements for the FID, KID, and CLIP cores. Yet, for the maximum number of views ($8$) in this experiments, we see that the FID and KID scores do not further improve. As shown in Figure 8, we observe more artifacts appearing in Paint3D, probably due to its use of mere backprojection which often leads to more seams when more views are backprojected. Our method scores with $8$ views are affected less; our neural backprojection produces smoother results, yet we do notice a bit more oversmoothing in our case, which is a limitation our our method.
Figure 8: Results for Paint3D and our method for $4$, $6$, and $8$ input views.  Im2SurfTex generates smoother surfaces.